OpenStudy (anonymous):

The spinner on a wheel of fortune can land with an equal chance on any one of ten regions. Three regions are red, four are blue, two are yellow, and one is green. A player wins $4 if the spinner stops on red and $2 if it stops on green. The player loses $1 if it stops on blue and $5 if it stops on yellow. What is the expected value rounded to the nearest cent?

OpenStudy (tkhunny):

What do you think? What is p(stops on red)? What is p(stops on green)? What is p(stops on blue)? What is p(stops on yellow)?

OpenStudy (anonymous):

p(4) p(2) p(1) p(5)

OpenStudy (tkhunny):

I think you might mean p(stops on red) = 4 p(stops on green) = 2 p(stops on blue) = 1 p(stops on yellow) = 5 Or something like that. Anyway, that's no good. Probabilities, if you have hte ALL, have to add up to one (1). You are given that there are ten (10) total, thus: p(stops on red) = 3/10 = 0.3 p(stops on green) = 1/10 = 0.1 p(stops on blue) = 4/10 = 0.4 p(stops on yellow) = 2/10 = 0.2 Add them up to check.

OpenStudy (anonymous):

ok thank you!